OpenStudy (anonymous):

If the domain of each of the functions f(x) and g(x) is all real numbers, will the domain of (f/g)(x) also be all real numbers? Why or why not?

OpenStudy (anonymous):

(f/g)(x)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(f/g)(x) = f(x)/g(x)

OpenStudy (anonymous):

hey u there?

Directrix (directrix):

I am voting "no" on this based on my reading here: http://math.ucsd.edu/~wgarner/math4c/textbook/chapter2/combfunctions.htm Suppose we have two functions: f(x) = sin(x) and g(x) = x². Both have domains that are the set of Real numbers. But, f(x)/g(x) = (f/g) (x) = sin(x) /x² . So, (f/g) (0) = sin(0)/0² which is undefined. Zero is not in the domain of (f/g) (x) although zero is in the domain of f(x) and of g(x)